OpenStudy (anonymous):

Eliminate the parameter t to find the cartesian equation of this curve. Find the points on the curve at which the tangent line is horizontal or vertical. Let x=2t-t^2 and y=-t+t^2

OpenStudy (anonymous):

Try solving for t on one of the equations.

OpenStudy (anonymous):

(Then substitute the answer into your other equation.)

OpenStudy (anonymous):

x=t(2-t) and y=t(t-1), can i plug in two different values for t, like t=y, t=y+1?

OpenStudy (cwrw238):